# Frequency-Explicit Shape Holomorphy in Uncertainty Quantification for Acoustic Scattering

### v1

@@ 1.1 Scattering Transmission Problem
In the acoustic modeling context, the solution $u$ of (1.1) gives the complex amplitude of the sound pressure of the so-called scattered wave, see [22, Sect 2.1] for more detail.
Excitation is provided by an incident wave $u^{\mathrm{inc}}$ that satisfies $-\Delta u^{\mathrm{inc}}-k^{2}u^{\mathrm{inc}}=0$ in $\mathbb{R}^{d}$, the prime example being a plane wave $u^{\mathrm{inc}}({\boldsymbol{x}})=\exp({\rm i}k{\boldsymbol{d}}\cdot{\boldsymbol{x}})$, propagating in direction ${\boldsymbol{d}}\in\mathbb{R}^{d}$, $|{\boldsymbol{d}}|=1$. Existence and uniqueness of a solution $u\in H^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$ of (1.1) is well-known, see, e.g., [53, Lemma 2.2].
@@ 1.2 Uncertainty Quantification (UQ) via Polynomial Surrogate Models
The shape of the scatterer $D$ is assumed to be uncertain, which we take into account by the customary approach to deterministic UQ through stochastic parameterization: the uncertainty in $D$ is captured by introducing a dependence of $D$ on (possibly infinitely many) real-valued random variables with known distributions. The key idea is to relegate those to mere real parameters. Thus we arrive at a family of transmission problems depending on infinitely many parameters. The solution $u$ of the transmission problem and any derived quantity of interest will then become functions of the parameters, alike.
Next we build sparse polynomial surrogate models of those functions on the parameter domain. This can be done accurately and efficiently using suitable spaces of multi-variate polynomials in the parameters, provided that $u$ and the quantities of interest are analytic/holomorphic [^2] We use “analytic” and “holomorphic” as synonyms, cf. [55, Definition 5.1]. functions of the parameters with a sufficiently large domain of analyticity in the complexified parameter space.
@@ 1.3 Simplest Case: Size Uncertainty Quantification
see [6, 7, 1]. To explain this growth, recall that resonances of the Helmholtz transmission problem are poles of the meromorphic extension $\widehat{k}\in\mathbb{C}\mapsto\widehat{u}(\widehat{k})$. When $n_{i}>1$ and $D_{0}$ is smooth with strictly-positive curvature there are resonances superalgebraically close to the real axis [57, Theorem 1.1], corresponding physically to “whispering-gallery modes” created by total internal reflection of rays hitting $\partial D_{0}$ from $D_{0}$. The $\widehat{k}_{j}$, often called quasi-resonances, can then be thought of as the real parts of these resonances close to the real axis; see Figure 1.1 (a) for a numerical illustration. We highlight also that the density of the near-real-axis resonances, and hence also the $\widehat{k_{j}}$, increases as $j$ increases [9, Theorem 1.3].
@@ 1.4 Scattering problems without quasi-resonances
We now explore the implications of the results from [53] in the simple size-UQ setting of Section 1.3. If $n_{i}\leq 1$ and $D_{0}$ is star-shaped with respect to 00, then $y\in\mathbb{C}\mapsto\widehat{u}(y)$ is meromorphic on $\left\{z\in\mathbb{C}:\,\operatorname{Re}z>{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{-2}}\right\}$ 2 }. However, since $\widehat{k}(y)=(1+\frac{1}{2}y)k$, the distance of the poles of $y\mapsto\widehat{u}(y)$ to the real axis shrinks like $O(k^{-1})$ as $k\to\infty$. This means that the extent of the domain of analyticity of the function $y\in\mathbb{C}\mapsto\widehat{u}(y)$ in the direction of the imaginary axis will decrease like $O(k^{-1})$ for $k\to\infty$. Therefore, even in non-resonant situations, $k$-uniformly accurate polynomial surrogate models for $y\mapsto\widehat{u}(y)$ will require increasingly larger degree for growing $k$. This growth with respect to $k$ can be avoided only if the size of admissible shape perturbations (controlled by the range of the parameter $y$ in the size-UQ model) decreases as $O(k^{-1})$ for $k\to\infty$. In other words,
@@
with $u$ satisfying the Sommerfeld radiation condition and $A({\boldsymbol{x}},{\boldsymbol{y}})$ and $n({\boldsymbol{x}},{\boldsymbol{y}})$ perturbations of some $A_{0}({\boldsymbol{x}})$ and $n_{0}({\boldsymbol{x}})$, respectively. When the problem with $A_{0}$ and $n_{0}$ is nontrapping, given $k_{0}>0$ there exists $C>0$ such that for all $k\geq k_{0}$ the map ${\boldsymbol{y}}\mapsto u({\boldsymbol{y}})$ is analytic for all ${\boldsymbol{y}}$ such that [61, Theorem 1.1]
@@
with this bound sharp through an unbounded sequence of wavenumbers $k$ [61, Theorem 1.4]. In [34, Theorem 4.2], the condition (1.10) is also shown to be a sufficient condition for analyticity (in the form of the relevant bounds on derivatives of $u$ with respect to ${\boldsymbol{y}}$) in the case when $A\equiv I$, the Sommerfeld radiation condition is approximated by an impedance boundary condition, and for particular classes of nontrapping $n_{0}$ and perturbations $n-n_{0}$.
@@ 1.5 Layout of the paper
In Section 3 we review the results of [53], apply them to the concrete transmission problem (1.1)-(1.2) and state more precisely the estimate (1.9) and generalizations of it, providing completely $k$-explicit expressions for the constants. Then, in Section 4, we admit $\mathbb{C}$-valued radial displacement functions $r$ in (1.11); thanks to the domain mapping approach the resulting transmission problem remains well defined and a perturbation argument yields $k$-explicit stability estimates. These estimates give detailed $k$-explicit information about shape holomorphy, more precisely, about the domain of analyticity of $r\mapsto\widehat{u}(r)$, $\widehat{u}(r)$ the solution of (1.1), with $D=D(r)$ given by (1.11), pulled back to the unit sphere. Such “shape holomorphy” results have been established in recent years for a host of parametric PDEs models [10, 2, 39, 40, 19, 45]. In these works, the size of the holomorphy domain of parametric solutions and operators was not made explicit in terms of problem parameters such as wavenumber, or (in [19]) the Reynolds number.
@@ 2.1 A parametric family of diffeomorphisms
As in [41, Section 3], we shift parameter dependence from the domain $D(r)$ to the coefficients of a transformed transmission problem, for which the scatterer occupies the unit ball $\widehat{D}:=\{{\boldsymbol{x}}\in\mathbb{R}^{d}:\,|{\boldsymbol{x}}|<1\}$. Taking the cue from [41, Equation (3.2)] or [35, Section 3.2], the transformation is effected by the parameter-dependent mapping $\boldsymbol{\Phi}(r):\mathbb{R}^{2}\to\mathbb{R}^{2}$, $r\in{\cal R}$, given by
@@
2.2 Mapped variational formulation of
@@
We repeat that the variational problem (2.7) depends on $r\in{\cal R}$ via $D=D(r)$ and (1.2). This dependence becomes more apparent in the transformed variational problem formed by pulling back (2.7) to the nominal setting with interface $\partial\widehat{D}=\mathbb{S}^{d-1}$ through the diffeomorphism $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$. As derived in [41, Section 4.1], this transformed variational problem reads:
@@
=$2\displaystyle=\boldsymbol{\Psi}({\boldsymbol{x}})+\begin{aligned} &\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{1}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{{\boldsymbol{x}}^{\top}{\boldsymbol{h}}}{|{\boldsymbol{x}}|}+\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{2}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{1}{|{\boldsymbol{x}}|}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{2}}}}\right){\boldsymbol{h}}\\ &+\frac{1}{|{\boldsymbol{x}}|}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{2}}\right){\boldsymbol{h}}+O(|{\boldsymbol{h}}|^{2}).\end{aligned}$
@@
The following result – equivalence of norms of functions under the transformation $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ – is a consequence of (2.16) and (2.17). For a proof, see [41, Lemma 3.4] and [58, Appendix E].
@@ 3 $k$-Explicit Norm Bounds for the Solution of the Transmission Problem
As discussed in §1.3-1.4, the behaviour of the norm of the solution operator of the transmission problem with respect to $k$ depends on whether $n_{i}<1$ or $n_{i}>1$; if $n_{i}<1$ the norm of the solution operator from $L^{2}(B_{2})\to L^{2}(B_{2})$ grows like $k$ (see (3.3) below); if $n_{i}>1$ the solution operator can grow exponentially through $k_{j}$ for a discrete sequence $0<k_{1}<k_{2}<\ldots\infty$ [57], [62], [53, §6], although the growth for “most” $k$ is at most algebraic in $k$ [46]. We now recall bounds from [53, Theorem 3.1] on the solution when $n_{i}<1$.
@@
The bound (3.3) is proved in [53, Theorem 3.1]; indeed, (3.3) follows from [53, Equation 3.2] by (i) choosing $a_{i}=a_{o}=A_{D}=A_{N}=n_{o}=1$ and (ii) by bounding $k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+n_{i}\|u\|_{L^{2}(B_{2})}^{2}$ below by $n_{i}(k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+\|u\|_{L^{2}(B_{2})}^{2})$, and (iii) noting that, in the proof of [53, Theorem 3.1], ${\rm diam}(\Omega_{i})$ can be replaced by $2$ (when going from [53, Equation 5.3] to [53, Equation 5.4]). Having proved the bound (3.3), the bound (3.5) follows by, e.g., [36, Theorem 5.1]. ∎
@@
The analogous $k$-explicit bounds on the solution of the Helmholtz exterior Dirichlet problem that hold uniformly for all Lipschitz star-shaped $D$ were proved in [11, Corollary 3.9] (see also [36, Theorem 2.5]). The appropriate analogues of the results in the next section about $k$-explicit holomorphic dependence of the solution on the radial displacement function could therefore also be obtained for the Helmholtz exterior Dirichlet problem using these bounds in place of Theorem 3.1.
@@ 4 $k$-Explicit Holomorphic Dependence of the Solution on the Radial Displacement Function
The $\mathfrak{r}$-dependent coefficients $\widehat{{\mathbf{A}}}(\mathfrak{r};\cdot)$ and $\widehat{n}(\mathfrak{r};\cdot)$ in the sesquilinear form on the left-hand side of (2.8)/(4.2) are holomorphic on $\mathfrak{R}$. Let $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))$ be defined as the left-hand side of (4.2) minus the right-hand side, so that $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))=0$. The analytic implicit function theorem, e.g. [23, Thm. 15.3], applied to $F$ and Part (i) of Theorem 4.1 then implies the following result.
@@
(Analyticity of $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})\in H^{2}$) By Part (ii) of Theorem 4.1, an analogous result holds for the mapping $\mathfrak{r}\in C^{2}(\mathbb{S}^{d-1},\mathbb{C})\mapsto\widehat{u}(\mathfrak{r})\in H^{2}(B_{2})$ when $\widehat{L}(\widehat{v}):=\int_{B_{2}}f\,\overline{\widehat{v}}$ for $f\in L^{2}(B_{2})$. Although we do not use this result in the rest of the paper, additional smoothness of the target space of the analytic map $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})$, translates into higher Sobolev regularity of the (analytic continuation of) the parametric solution. The higher Sobolev regularity of the analytic continuation of the parametric solution (as opposed to mere boundedness of this continuation in the energy norm of the PDE) is key in the convergence analysis of multilevel QMC methods; see, e.g., [24, §4.3.1] and the further discussion in §9.
@@
By, e.g., [50, Page 122], the condition (4.8) ensures that the differential operator underlying the variational formulation (4.2) is strongly elliptic. We note that (4.8) is analogous to the condition [41, Equation 5.16] in [41, Assumption 5.11].
@@
By Corollary 4.6, the sesquilinear form $\widehat{a}(\cdot,\cdot)$ is continuous and satisfies a Gårding inequality, with constants independent of $\mathfrak{r}$. Therefore, by Fredholm theory, if, under the assumption of existence, one has a bound on the solution of the Helmholtz transmission problem in terms of the data $\widehat{L}$, then the solution exists and is unique (for more details of this method of arguing applied to the Helmholtz equation, see, e.g., [36, Lemma 3.5]).
@@
Then, by the $H^{2}$ regularity result of [47, Theorem 6.1], there exists $C>0$ (depending on $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{\boldsymbol{\nabla}}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, and the constant $c_{2}$ in (4.8)) such that, for all $k\geq 1/2$, say,
@@
Given $z_{1},z_{2}\in\mathbb{C}$ with $|z_{1}|,|z_{2}|\geq L_{0}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{>}}0$, for any two elements ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in V$ of a normed vector space $V$,
@@
5 FE Discretization: $k$-Explicit Error Bounds
In §4 we proved existence, uniqueness, and $k$-explicit bounds on the solution $\widehat{u}(\mathfrak{r};\cdot)$ to the variational problem (4.2), both for general data in $(H_{k}^{1}(B_{2}))^{*}$ and for $L^{2}$ data, see Theorem 4.1.
In this section we focus on computing approximations to $\widehat{u}(\mathfrak{r};\cdot)$, considering both general data in $(H_{k}^{1}(B_{2}))^{*}$ and the special case of $L^{2}$ data; the relevance of the latter is that the plane-wave scattering problem (1.1) can be reformulated as the solution of a problem with $L^{2}$ data – see §6.2 below.
Since the Dirichlet-to-Neumann map $\mathsf{DtN}$ is expensive to compute, we follow the commonly-used practice of approximation the radiation condition by truncation of the unbounded exterior domain by a radial perfectly-matched layer (PML).

### v2

@@ 1.1 Scattering Transmission Problem
In the acoustic modeling context, the solution $u$ of (1.1) gives the complex amplitude of the sound pressure of the so-called scattered wave, see [23, Sect 2.1] for more detail.
Excitation is provided by an incident wave $u^{\mathrm{inc}}$ that satisfies $-\Delta u^{\mathrm{inc}}-k^{2}u^{\mathrm{inc}}=0$ in $\mathbb{R}^{d}$, the prime example being a plane wave $u^{\mathrm{inc}}({\boldsymbol{x}})=\exp({\rm i}k{\boldsymbol{d}}\cdot{\boldsymbol{x}})$, propagating in direction ${\boldsymbol{d}}\in\mathbb{R}^{d}$, $|{\boldsymbol{d}}|=1$. Existence and uniqueness of a solution $u\in H^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$ of (1.1) is well-known, see, e.g., [56, Lemma 2.2].
@@ 1.2 Uncertainty Quantification (UQ) via Polynomial Surrogate Models
The shape of the scatterer $D$ is assumed to be uncertain, which we take into account by the customary approach to deterministic UQ through stochastic parameterization: the uncertainty in $D$ is captured by introducing a dependence of $D$ on (possibly infinitely many) real-valued random variables with known distributions, see, e.g., [17] for this approach and examples. The key idea is to relegate those to real-valued, deterministic parameters, and to endow the (possibly infinite, but countable) cartesian product of the parameter domains with probability measures. We refer to [17, 16, 47, 2] and the references therein for discussions and realizations of this idea. In the present paper, we arrive at a family of transmission problems depending on infinitely many parameters. Holomorphic dependence of solutions on the shape of the domain in which the PDE is set is available for a wide range of elliptic and parabolic PDEs, see [43, 47, 20, 42, 2] and the references there. The results in these references were not explict in e.g. the wavenumber. Frequency-explicit holomorphy of solutions of time-harmonic, acoustic scattering was developed for certain forward and problems acoustic scattering problems recently in [24, 48, 38, 36]. The solution $u$ of the presently considered transmission problem and any derived quantity of interest will then become functions of the parameters, alike.
Next we build sparse polynomial surrogate models of those functions on the parameter domain. This can be done accurately and efficiently using suitable spaces of multi-variate polynomials in the parameters, provided that $u$ and the quantities of interest are analytic/holomorphic [^2] We use “analytic” and “holomorphic” as synonyms, cf. [58, Definition 5.1]. functions of the parameters with a sufficiently large domain of analyticity in the extension of the parameter space into the complex domain; see [17] and the works cited there.
@@ 1.3 Simplest Case: Size Uncertainty Quantification
see [6, 7, 1]. To explain this growth, recall that resonances of the Helmholtz transmission problem are poles of the meromorphic extension $\widehat{k}\in\mathbb{C}\mapsto\widehat{u}(\widehat{k})$. When $n_{i}>1$ and $D_{0}$ is smooth with strictly-positive curvature there are resonances superalgebraically close to the real axis [60, Theorem 1.1], corresponding physically to “whispering-gallery modes” created by total internal reflection of rays hitting $\partial D_{0}$ from $D_{0}$. The $\widehat{k}_{j}$, often called quasi-resonances, can then be thought of as the real parts of these resonances close to the real axis; see Figure 1.1 (a) for a numerical illustration. We highlight also that the density of the near-real-axis resonances, and hence also the $\widehat{k_{j}}$, increases as $j$ increases [9, Theorem 1.3].
@@ 1.4 Scattering problems without quasi-resonances
We now explore the implications of the results from [56] in the simple size-UQ setting of Section 1.3. If $n_{i}\leq 1$ and $D_{0}$ is star-shaped with respect to 00, then $y\in\mathbb{C}\mapsto\widehat{u}(y)$ is meromorphic on $\left\{z\in\mathbb{C}:\,\operatorname{Re}z>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2}\right\}$. However, since $\widehat{k}(y)=(1+\frac{1}{2}y)k$, the distance of the poles of $y\mapsto\widehat{u}(y)$ to the real axis shrinks like $O(k^{-1})$ as $k\to\infty$. This means that the extent of the domain of analyticity of the function $y\in\mathbb{C}\mapsto\widehat{u}(y)$ in the direction of the imaginary axis will decrease like $O(k^{-1})$ for $k\to\infty$. Therefore, even in non-resonant situations, $k$-uniformly accurate polynomial surrogate models for $y\mapsto\widehat{u}(y)$ will require increasingly larger degree for growing $k$. This growth with respect to $k$ can be avoided only if the size of admissible shape perturbations (controlled by the range of the parameter $y$ in the size-UQ model) decreases as $O(k^{-1})$ for $k\to\infty$. In other words,
@@
with $u$ satisfying the Sommerfeld radiation condition and $A({\boldsymbol{x}},{\boldsymbol{y}})$ and $n({\boldsymbol{x}},{\boldsymbol{y}})$ perturbations of some $A_{0}({\boldsymbol{x}})$ and $n_{0}({\boldsymbol{x}})$, respectively. When the problem with $A_{0}$ and $n_{0}$ is nontrapping, given $k_{0}>0$ there exists $C>0$ such that for all $k\geq k_{0}$ the map ${\boldsymbol{y}}\mapsto u({\boldsymbol{y}})$ is analytic for all ${\boldsymbol{y}}$ such that [64, Theorem 1.1]
@@
with this bound sharp through an unbounded sequence of wavenumbers $k$ [64, Theorem 1.4]. In [36, Theorem 4.2] and [38, Theorem 3.1], the condition (1.10) is also shown to be a sufficient condition for analyticity (in the form of the relevant bounds on derivatives of $u$ with respect to ${\boldsymbol{y}}$) for particular classes of nontrapping $A_{0},n_{0}$. In [48, Section 4] $O(k^{-1})$ bounds on shape perturbations are also identified as necessary for reliable Bayesian shape inversion.
@@ 1.5 Layout of the paper
In Section 3 we review the results of [56], apply them to the concrete transmission problem (1.1)-(1.2) and state more precisely the estimate (1.9) and generalizations of it, providing completely $k$-explicit expressions for the constants. Then, in Section 4, we admit $\mathbb{C}$-valued radial displacement functions $r$ in (1.11); thanks to the domain mapping approach the resulting transmission problem remains well defined and a perturbation argument yields $k$-explicit stability estimates. These estimates give detailed $k$-explicit information about shape holomorphy, more precisely, about the domain of analyticity of $r\mapsto\widehat{u}(r)$, $\widehat{u}(r)$ the solution of (1.1), with $D=D(r)$ given by (1.11), pulled back to the unit sphere. Such “shape holomorphy” results have been established in recent years for a host of parametric PDEs models [10, 2, 42, 20, 47]. In these works, the size of the holomorphy domain of parametric solutions and operators was not made explicit in terms of problem parameters such as wavenumber, or (in [20]) the Reynolds number.
@@ 2.1 A parametric family of diffeomorphisms
As in [43, Section 3], we shift parameter dependence from the domain $D(r)$ to the coefficients of a transformed transmission problem, for which the scatterer occupies the unit ball $\widehat{D}:=\{{\boldsymbol{x}}\in\mathbb{R}^{d}:\,|{\boldsymbol{x}}|<1\}$. Taking the cue from [43, Equation (3.2)] or [37, Section 3.2], the transformation is effected by the parameter-dependent mapping $\boldsymbol{\Phi}(r):\mathbb{R}^{2}\to\mathbb{R}^{2}$, $r\in{\cal R}$, given by
@@
2.2 Variational formulation under pullback via the domain mapping $\Phi$
@@
We repeat that the variational problem (2.7) depends on $r\in{\cal R}$ via $D=D(r)$ and (1.2). This dependence becomes more apparent in the transformed variational problem formed by pulling back (2.7) to the nominal setting with interface $\partial\widehat{D}=\mathbb{S}^{d-1}$ through the diffeomorphism $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$. As derived in [43, Section 4.1], this transformed variational problem reads:
@@
=$2\displaystyle=\boldsymbol{\Psi}({\boldsymbol{x}})+\begin{aligned} &\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{1}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{{\boldsymbol{x}}^{\top}{\boldsymbol{h}}}{|{\boldsymbol{x}}|}+\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{2}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{1}{|{\boldsymbol{x}}|}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{2}}}}\right){\boldsymbol{h}}\\ &+\frac{1}{|{\boldsymbol{x}}|}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{2}}\right){\boldsymbol{h}}+O(|{\boldsymbol{h}}|^{2}).\end{aligned}$
@@
The following result – equivalence of norms of functions under the transformation $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ – is a consequence of (2.16) and (2.17). For a proof, see [43, Lemma 3.4] and [61, Appendix E].
@@ 3 $k$-Explicit Norm Bounds for the Solution of the Transmission Problem
As discussed in §1.3-1.4, the behaviour of the norm of the solution operator of the transmission problem with respect to $k$ depends on whether $n_{i}<1$ or $n_{i}>1$; if $n_{i}<1$ the norm of the solution operator from $L^{2}(B_{2})\to L^{2}(B_{2})$ grows like $k$ (see (3.3) below); if $n_{i}>1$ the solution operator can grow exponentially through $k_{j}$ for a discrete sequence $0<k_{1}<k_{2}<\ldots\infty$ [60], [65], [56, §6], although the growth for “most” $k$ is at most algebraic in $k$ [49]. We now recall bounds from [56, Theorem 3.1] on the solution when $n_{i}<1$.
@@
The bound (3.3) is proved in [56, Theorem 3.1]; indeed, (3.3) follows from [56, Equation 3.2] by (i) choosing $a_{i}=a_{o}=A_{D}=A_{N}=n_{o}=1$ and (ii) by bounding $k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+n_{i}\|u\|_{L^{2}(B_{2})}^{2}$ below by $n_{i}(k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+\|u\|_{L^{2}(B_{2})}^{2})$, and (iii) noting that, in the proof of [56, Theorem 3.1], ${\rm diam}(\Omega_{i})$ can be replaced by $2$ (when going from [56, Equation 5.3] to [56, Equation 5.4]). Having proved the bound (3.3), the bound (3.5) follows by, e.g., [39, Theorem 5.1]. ∎
@@
The analogous $k$-explicit bounds on the solution of the Helmholtz exterior Dirichlet problem that hold uniformly for all Lipschitz star-shaped $D$ were proved in [11, Corollary 3.9] (see also [39, Theorem 2.5]). The appropriate analogues of the results in the next section about $k$-explicit holomorphic dependence of the solution on the radial displacement function could therefore also be obtained for the Helmholtz exterior Dirichlet problem using these bounds in place of Theorem 3.1.
@@ 4 $k$-Explicit Holomorphic Dependence of the Solution on the Radial Displacement Function
The $\mathfrak{r}$-dependent coefficients $\widehat{{\mathbf{A}}}(\mathfrak{r};\cdot)$ and $\widehat{n}(\mathfrak{r};\cdot)$ in the sesquilinear form on the left-hand side of (2.8)/(4.2) are holomorphic on $\mathfrak{R}$. Let $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))$ be defined as the left-hand side of (4.2) minus the right-hand side, so that $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))=0$. The analytic implicit function theorem, e.g. [25, Theorem 15.3], applied to $F$ and Part (i) of Theorem 4.1 then implies the following result.
@@
(Analyticity of $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})\in H^{2}$) By Part (ii) of Theorem 4.1, an analogous result holds for the mapping $\mathfrak{r}\in C^{2}(\mathbb{S}^{d-1},\mathbb{C})\mapsto\widehat{u}(\mathfrak{r})\in H^{2}(B_{2})$ when $\widehat{L}(\widehat{v}):=\int_{B_{2}}f\,\overline{\widehat{v}}$ for $f\in L^{2}(B_{2})$. Although we do not use this result in the rest of the paper, additional smoothness of the target space of the analytic map $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})$, translates into higher Sobolev regularity of the (analytic continuation of) the parametric solution. The higher Sobolev regularity of the analytic continuation of the parametric solution (as opposed to mere boundedness of this continuation in the energy norm of the PDE) is key in the convergence analysis of multilevel QMC methods; see, e.g., [26, §4.3.1] and the further discussion in §9.
@@
By, e.g., [53, Page 122], the condition (4.8) ensures that the differential operator underlying the variational formulation (4.2) is strongly elliptic. We note that (4.8) is analogous to the condition [43, Equation 5.16] in [43, Assumption 5.11].
@@
By Corollary 4.6, the sesquilinear form $\widehat{a}(\cdot,\cdot)$ is continuous and satisfies a Gårding inequality, with constants independent of $\mathfrak{r}$. Therefore, by Fredholm theory, if, under the assumption of existence, one has a bound on the solution of the Helmholtz transmission problem in terms of the data $\widehat{L}$, then the solution exists and is unique (for more details of this method of arguing applied to the Helmholtz equation, see, e.g., [39, Lemma 3.5]).
@@
Then, by the $H^{2}$ regularity result of [50, Theorem 6.1], there exists $C>0$ (depending on $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{\boldsymbol{\nabla}}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, and the constant $c_{2}$ in (4.8)) such that, for all $k\geq 1/2$, say,
@@
Given $z_{1},z_{2}\in\mathbb{C}$ with $|z_{1}|,|z_{2}|\geq L_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{>}}0$, for any two elements ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in V$ of a normed vector space $V$,
@@
5 Recap of $k$-explicit Finite-Element Error Bounds
In §4 we proved existence, uniqueness, and $k$-explicit bounds on the solution $\widehat{u}(\mathfrak{r};\cdot)$ to the variational problem (4.2); see Theorem 4.1. In this section we recall results from [33, 35] about computing approximations to $\widehat{u}(\mathfrak{r};\cdot)$ using the $h$-version of the FEM (i.e., where the polynomial degree is fixed and accuracy is increased by decreasing $h$) with the radiation condition approximated by a radial perfectly-matched layer (PML). When applied to our set up, [33, 35] establish FEM error bounds for each $\mathfrak{r}\in\mathcal{A}$ (4.7), with the constants in principle depending on $\mathfrak{r}$. In this section we go through some of the arguments in [33, 35] to justify that the constants in these error bounds can be taken to be independent of $\mathfrak{r}\in\mathcal{A}$.